OpenStudy (anonymous):

What is the slope intercept form of the equation x - 3y= 3?

OpenStudy (anonymous):

y = -1/3x - 2??

OpenStudy (gorv):

@InsanelyChaotic y=x/3-1

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

y =1/3 x - 1

OpenStudy (gorv):

@InsanelyChaotic it is right i m 100% sure

OpenStudy (gorv):

slope intercept form is y=mx+c convert ur euation in this form u will get the answer